OpenStudy (anonymous):

Where does the graph of y = 6x2 + 5x − 6 cross the x-axis? (−1/6, 0) and (6, 0) (1/6, 0) and (−6, 0) (2/3, 0) and (−, 0) (−2/3, 0) and (, 0)

hero (hero):

1. Set y = 0: 6x^2 + 5x - 6 = 0 2. Split the middle term: 6x^2 + 9x - 4x - 6 = 0 3. Factor the first two terms, then factor the last two terms to get: 3x(2x + 3) - 2(2x + 3) = 0 4. Since 2x + 3 is common to both terms, factor it out to get: (2x + 3)(3x - 2) = 0 Finish solving for x to find the roots

OpenStudy (anonymous):

how do i solve for x

OpenStudy (anonymous):

x=-3/2 -1 x=

OpenStudy (anonymous):

x = -1 srry

hero (hero):

Sorry, I meant to say to solve 2x + 3 = 0 3x - 2 = 0

OpenStudy (anonymous):

x= -3/2 x=2/3